Fermi National Laboratory


Accelerator Update

Wednesday August 18
The day shift began with Operations monitoring store 3739, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

Operations established beam to the Meson MTest and MCenter experiments at 9:02 AM.

At 10 AM, Operations halted stacking so that Booster experts could load new code for a power supply (GMPS), so that a Pbar tech could replace a kicker (D:IKIK) module, and so a Main Injector (MI) RF expert could work on a couple of RF stations. Stacking resumed at 10:24 AM.

Pbar experts began a reverse proton study and a MI expert began conducting aperture scans at 1 PM.

Operations reported at 2:25 PM that the Neutron Therapy Facility (NTF) began, over the next forty minutes, taking occasional spurts of beam.

At 6:28 PM, a Booster RF station (BRF5) tripped off for the first of four times in four minutes. Two Booster vacuum pumps tripped off due to the RF station trips. Everything reset okay.

Recycler experts began a short study at 6:55 PM and ended around fifteen minutes later.

Pbar studies ended at 8:46 PM.

Pbar resumed stacking at 9:07 PM.

The MI expert finished his scans at midnight.

Thursday August 19
The midnight shift began with Operations monitoring store 3739, with Pbar stacking, and with the MiniBooNE, MTest, and MCenter experiments taking beam.

A MI expert began a damper study at 1:11 AM and finished around 2:30 AM.

Operations terminated store 3739 at 9:01 AM.

Operations prepared the Collider Detector at Fermilab (CDF) and D-Zero collision halls for access at 9:14 AM.

Operations set up the Switchyard (SY) beam to go only to the SY dump for a study at 9:35 AM; there was no MCenter or MTest beam during the study.

Both CDF and D0 were out of access by 10:51 AM.

A Pbar expert began an AP1 transfer line study at 1:44 PM.

NTF took about thirty minutes of beam for studies beginning at 2:35 PM.

At 3:39 PM, the Tevatron (TeV) lost its ramp. The TeV sector A1 cryo system went bad, but this wasn't the cause of the ramp problem. Investigations continued.

At 3:48 PM, all the MI RF stations tripped off for no apparent reason. Everything reset okay.

TeV experts reported at 4 PM that there was a TeV ground fault.

At 4:26 PM, three MI RF stations (MIRF7, 17 & 18) tripped off a bunch of times over the next hour. Operators reset the stations after each trip.

TeV experts reported at 4:47 PM that the TeV ground fault was somewhere between TeV sectors F2 and A2. Operations prepared the TeV tunnels for access. This access halted all SY beam.

TeV experts reported at 5:40 PM that they needed to access F-sector. Operators prepared the area. This access halted beam to Pbar and MiniBooNE.

TeV experts reported at 7:36 PM that they found a corroded voltage tap on a sector F35 spool piece and a pin-hole LCW leak in F4, but that neither of these things were the cause of the ground fault.

Friday August 20
The midnight shift began with the TeV ground fault search continuing, and with Operations monitoring the Pbar stack.

At 6:14 AM, Operations called in more experts to help with the TeV ground fault search.

Plans for Friday and the Weekend
The plan for today is to fix the TeV and establish a store. On Monday, the Fermi accelerator complex begins its scheduled thirteen-week shutdown. Hopefully, the TeV ground fault will be found and fixed, but if it isn't the shutdown schedule is not likely to change. There are studies that the machine experts need to perform before the shutdown begins.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 8/20/2004   email Fermilab